### v1

# Accelerated Mirror Descent for Non-Euclidean Star-convex Functions

Abstract
Acceleration for non-convex functions has been an important problem in optimisation. We revisit star-convex functions, which are strictly unimodal on all lines through a minimizer. In [18], the authors accelerate gradient descent for star-convex functions with gradients that are Lipschitz with respect to the Euclidean norm in an unconstrained domain. In this paper, we introduce a new assumption about the regularity of the derivative of a general norm and we accelerate mirror descent for this class of normed spaces. We show that, under it, our algorithms show sharp convergence rates for star-convex functions with $\alpha$-Hölder continuous gradients. We also prove that our convergence rate is near optimal for $p$-norms.
1 Introduction
Nesterov’s accelerated gradient descent [26] has proven to be a powerful tool for first-order optimization, significantly improving algorithmic performance. Nowadays several extensions of accelerated methods exist including coordinate descent methods [27, 12], distributed gradient descent [30] and proximal methods [13] with applications across widespread domains such as image deblurring [2], compressed sensing [3] and deep learning [31].
Theoretical results analysing accelerated methods mainly focus on smooth convex optimisation, demonstrating optimal convergence and improved performance against traditional gradient methods, see [26] and [6] for further references. However, this framework becomes less applicable to modern machine learning applications due to their typically non-convex energy landscapes. Attempts to understand acceleration for non-convex functions has been made, for instance, by leveraging the Łojasiewicz inequalities [5]. Further classes of non-convex functions are discussed in [7] and [18](Appendix A.1).
[28] relaxes the convexity assumption to so-called star-convex functions, a class of "structured" non-convex function, which are strictly unimodular on all lines through a minimizer. As convex functions exhibit unimodularity along all lines, star-convexity represents a weaker condition. [23] provide examples of functions that are star-convex but not convex. Additionally, [36] and [22] provide evidence that neural network loss functions often satisfy the star-convexity (but not the convexity) assumption, emphasizing its relevance for modern machine learning applications.
For star-convex functions, [18] present an accelerated framework using binary search techniques that is nearly optimal provided that the objective function has a Lipschitz gradient in Euclidean norm. However, recent machine learning results such as [32, 1] minimize the objective functions in non-Euclidean norms, which are not covered by the findings in [18]. This gap inspires us to extend and generalize their algorithm for arbitrary normed spaces. We identify a new regularity condition for norms (Definition 2.7) that enables acceleration. In regular normed spaces, we combine a generalized "line-search" style algorithm [18] (Algo 2) with accelerated mirror descent to minimize star-convex functions with $\alpha$-Hölder continuous gradients (also known as weakly smooth functions, Definition 2.4). This indicates that compensating for the non-convexity in the objective function requires regularity in the norm rather than the function itself. Our finding challenges the common belief that the properties of the objective function are more crucial than those of the norm.
We demonstrate that in case of $p$-norms, which are regular for $p>1$ (see Lemma 2.1), our algorithm is nearly optimal. By showing that non-convex functions can be solved by accelerated algorithms in non-Euclidean geometry, we extend the traditional understanding of which functions admit acceleration. Our work suggests that more non-convex functions can be solved efficiently, potentially sparking increased interest in this area of research.
[…]
1.1 Literature review
Previous results on star-convex functions primarily focus on non-accelerated algorithms. [14] and [17] show convergence guarantees for stochastic gradient descent, [28] analyse cubic regularization and [23] investigate a cutting plane method. In [20], the star-convexity assumption is applied to stochastic and online learning settings, showing regret bounds for first-order methods, leading to a $\mathcal{O}(1/T)$ convergence rate (with $T$ being the number of gradient steps) in case that either the objective function or its gradient is Lipschitz.
Acceleration on star-convex functions has been studied in, e.g., [15] for low-dimensional subspace optimization and in [35] for an one-dimensional line search-type of algorithm, both showing convergence rates to an $\epsilon$-optimal point. However, as noted in [18], these algorithms can be limited by their potentially high computational costs for minimizing general star-convex functions.
[18] proposes an accelerated framework using binary search, showing near-optimal performance, provided the $\tau$-star-convex functions (see Definition 2.1) are $L$-smooth (see Definition 2.3) with respect to the Euclidean norm in a domain bounded by $R$. Similar to [15], their algorithm finds an $\epsilon$-optimal solution in $\mathcal{O}(\epsilon^{-1/2}L^{1/2}\tau R\log(LR^{2}/\epsilon))$ iterations. For smooth star-convex functions, they also provide a lower bound of $\Omega(L^{1/2}\tau R\epsilon^{-1/2})$.
While the accelerated framework in [18] is limited to smooth star-convex functions with respect to the Euclidean norm, Nesterov’s accelerated framework ([11, 24]) applies to general normed spaces under a weak smoothness (see Definition 2.4) assumption.
1.2 Our contributions
In this paper, we generalize the accelerated framework introduced in [18] to encompass weakly smooth star-convex functions with respect to a general norm. Addressing the challenges arising from the non-convexity of the objective function and the non-Euclidean geometry of the norm, we introduce new concepts, such as regular norms, and develop novel proof strategies for the algorithms convergence. Our key contributions are summarized as follows:
Acceleration on regular norms. To extend the acceleration framework of [18] from Euclidean norms to general norms, we identify sufficient conditions on the derivative of the squared norm $\|\cdot\|^{2}$. Introducing the notion of regular norms, we demonstrate that star-convexity and weak smoothness suffice to apply our generalized accelerated framework. Our results underscore the critical role of norm regularity, rather than the function properties, in achieving acceleration, encouraging further research in this direction.
Extended binary search. We generalize the binary search algorithm introduced in [18], originally designed for smooth functions to encompass weakly smooth objective functions. Additionally, we present a streamlined convergence proof (Thm 4.1) for our generalized algorithm.
Nearly optimal performance for $p$-norm. For $p$-norms, we extend the convergence rates in [24] from convex to star-convex functions. According to the lower bounds in [16], our rates are nearly optimal, up to some logarithmic factor. Table 1 categorises our results in the existing literature.
$p$ Bound Convex $\tau$-star-convex
$(1,2]$ Upper $\mathcal{O}_{\kappa}\left(\left(\frac{LR^{\kappa}}{\epsilon}\right)^{\frac{2}{3\kappa-2}}\right)$ $\mathcal{O}_{\kappa}\left(\left(\frac{L\tau^{2}R^{\kappa}}{\epsilon}\right)^{\frac{2}{3\kappa-2}}\log\Big{(}\frac{LR^{\kappa}}{\epsilon}\Big{)}\right)$
[24] Our work (Corollary 3.3)
Lower $\Omega_{\kappa}\left(\left(\frac{LR^{\kappa}}{\epsilon[\ln d]^{\kappa-1}}\right)^{\frac{2}{3\kappa-2}}\right)$
[16]
$[2,\infty]$ Upper $\mathcal{O}_{p,\kappa}\left(\left(\frac{LR^{\kappa}}{\epsilon}\right)^{\frac{p}{\kappa p+\kappa-p}}\right)$ $\mathcal{O}_{p,\kappa}\left(\left(\frac{L\tau^{p}R^{\kappa}}{\epsilon}\right)^{\frac{p}{\kappa p+\kappa-p}}\log\Big{(}\frac{LR^{\kappa}}{\epsilon}\Big{)}\right)$
[24] Our work (Corollary 3.4)
Lower $\Omega_{p,\kappa}\left(\left(\frac{LR^{\kappa}}{\epsilon}\right)^{\frac{p}{\kappa p+\kappa-p}}\right)$
[16]
Table 1: Summary of the relation between related existing work and our work for $\tau$-star-convex functions. $R$ is the radius of the domain of the $\tau$-star-convex objective function, $(L,\kappa)$ is its smoothness with respect to a $p$-norm $\|\cdot\|_{p}$ with $p>1$ and $\epsilon$ the accuracy of the algorithm. Since convex functions are $1$-star-convex, the lower bounds from [16] apply to both.
Paper outline. In Section 2, we outline the problem setting and provide preliminaries, including the norm regularity assumption and its properties. Section 3 presents our generalized accelerated algorithm and discusses its convergence guarantees for various settings. Section 4 introduces our binary search algorithm and its running time, and Section 5 concludes with a discussion on future research directions. All proofs are provided in the Appendix.
2 Problem setting and preliminaries
In this paper we consider the unconstrained minimization problem
$\min_{x\in\mathbb{R}^{d}}F(x)$
[…]
2.1 Definitions
In this paper, we consider the class of $\tau$-star-convex functions, as introduced in [17], which are defined as follows.
Definition 2.1 ($\tau$-star-convexity).
A continuously differentiable function $F:\mathcal{X}\xrightarrow{}\mathbb{R}$ is said to be $\tau$-star-convex for $\tau>0$ if there exists $x_{\star}\in\mathcal{X}$ such that for all $x\in\mathcal{X}$
$\tau\langle\nabla F(x),x-x_{\star}\rangle\geq F(x)-F(x_{\star}).$ (1)
Note that, all convex functions are $1$-star-convex [20]. If included, 00-star convexity is trivial.
[…]
Definition 2.2 (Bregman divergence).
For a continuously differentiable function $F:\mathcal{X}\xrightarrow{}\mathbb{R}$, the Bregman divergence is defined as
$D_{F}(x,y):=F(x)-F(y)-\langle\nabla F(y),x-y\rangle.$
Definition 2.3 (Smoothness).
[…]
$|D_{F}(x,y)|\leq\frac{L}{2}\|x-y\|^{2}.$
Definition 2.4 (Weak smoothness).
[…]
$|D_{F}(x,y)|\leq\frac{L}{\kappa}\|x-y\|^{\kappa}.$
When $\kappa=2$, weak smoothness becomes the usual smoothness. In the general case, we would refer to $\kappa<2$.
To develop our mirror descent algorithm, we use a distance-generating function (see Assumption 1), that fulfills strong or uniform convexity assumptions with respect to a norm, defined as follows.
Definition 2.5 (Strong convexity).
[…]
$D_{F}(x,y)\geq\frac{\mu}{2}\|x-y\|^{2}.$
Definition 2.6 (Uniform convexity).
[…]
$D_{F}(x,y)\geq\frac{\mu}{q}\|x-y\|^{q}.$
When $q=2$, it is equivalent to the usual strong convexity. In the general case, we would have $q>2$.
2.2 Norm regularity assumption
Without convexity of the objective function, acceleration is possible with additional assumptions, usually by constraining the class of objective functions. In our paper, we introduce a new regularity property for norms that we will show is a sufficient condition for acceleration.
[…]
$\phi(x):=\nabla\left(\frac{\|x\|^{2}}{2}\right).$
Definition 2.7 (Regular norm).
[…]
$\begin{cases}\langle\phi(x),x\rangle&\geq\|x\|^{2}\\ \|\phi(x)\|_{\ast}&\leq C\|x\|.\end{cases}$
Given that our setting is finite dimensional, the equivalence of all norms suggests a potentially large family of regular norms. However, to provide dimension-independent performance guarantees for our algorithm, we require the norms to be $C$-regular, with $C>0$ being independent of the dimension.
Since $\phi$ is a homogeneous function by definition, the regularity can be rewritten as
$\begin{cases}\inf_{y\in\mathbb{U}}\quad\langle\phi(y),y\rangle\geq 1\\ \sup_{y\in\mathbb{U}}\quad\|\phi(y)\|_{\ast}\leq C,\\ \end{cases}$
where $\mathbb{U}$ denotes the unit sphere. The first condition ensures compatibility of the norm with the scalar product, while the second condition requires the norm’s derivative operator to be bounded. These two properties are crucial in our accelerated framework (Algorithm 1), especially for weakly smooth objective functions. They allow for the choice of aggressive step-sizes and the derivation of accelerated convergence rates.
The following lemmas provide examples of norms that are $C$-regular with constant $C>0$ not depending on the input dimension $d$. The proofs can be found in Appendix A.
Lemma 2.1.
For $p>1$, the $p$-norm $\|\cdot\|_{p}$ is $1$-regular. The matrix Schatten $p$-norm is also $1$-regular.
Another example of a regular norm is the following: Fix $\lambda\in[0,1]$, and consider a norm $\|\cdot\|_{a}$ defined on $\mathbb{R}^{d_{a}}$ and a norm $\|\cdot\|_{b}$ defined on $\mathbb{R}^{d_{b}}$. The composite norm $\|\cdot\|_{a\circ b}$ is defined on $\mathbb{R}^{d_{a}+d_{b}}$ by
$\|x\|_{a\circ b}=\left\|(x_{a},x_{b})^{T}\right\|_{a\circ b}:=\sqrt{\lambda\|x_{a}\|_{a}^{2}+(1-\lambda)\|x_{b}\|_{b}^{2}},$
where $x_{a}:=x_{1:d_{a}}$ and $x_{b}:=x_{(d_{a}+1):(d_{a}+d_{b})}$.
Lemma 2.2.
Let norm $\|\cdot\|_{a}$ be $C_{a}$-regular and norm $\|\cdot\|_{b}$ be $C_{b}$-regular, then the composite norm $\|\cdot\|_{a\circ b}$ is $\max\{C_{a},C_{b}\}$-regular.
Lemma 2.2 indicates that our algorithm provides convergence guarantees not only for $p$-norms, but also for compositions of regular norms. Combined with Lemma 2.1, this shows that our algorithm is applicable for a wide range of non-Euclidean norms. For instance, since $\|\cdot\|_{2}$ and $\|\cdot\|_{1.5}$ are $1$-regular according to Lemma 2.1, Lemma 2.2 implies that the following composite norm is is also $1$-regular
$\|x\|_{2\circ 1.5}:=\sqrt{0.5\|x_{1:d/2}\|^{2}_{2}+0.5\|x_{(d/2+1):d}\|^{2}_{1.5}}.$ (2)
3 Main results
In this section, we develop our accelerated optimization algorithm for weakly smooth star-convex functions. We start by summarizing our assumptions.
Assumption 1.
[…]
  • $F$ is $\tau$-star-convex (Definition 2.1).
[…]
Our method is given in Algorithm 1. It extends the accelerated method of [18], which addresses the case of smoothness ($\kappa=2$) for the Euclidean norm $\left\lVert\cdot\right\rVert_{2}$ known to be $C=1$-regular and strongly convex with respect to itself $(q=2,\psi(\cdot)=\frac{1}{2}\left\lVert\cdot\right\rVert_{2}^{2})$. Similar to their method, ours maintains an auxiliary sequence $(x_{t})_{t\geq 1}$ of iterates, from which it derives a sequence $(x_{t}^{md})_{t\geq 1}$ of gradient query points as well as the output sequence $(x_{t}^{ag})_{t\geq 1}$ of approximate minimizers. In each iteration, the gradient point $x_{t}^{md}$ is found by binary search, which we extend from [18] to handle weak smoothness in Section 4. While this search is not necessary for convex $F$, it appears to be a reasonable trade-off for the relaxation to star-convexity, see also the discussion in Section 5. The iterates $x_{t}$ and approximate minimizers $x_{t}^{ag}$ are each updated in [18] using gradient descent. However, to address a non-Euclidean norm, we instead employ a mirror descent step with respect to $D^{\psi}$ to update $x_{t}$, and a proximal step with respect to $\frac{1}{q}\left\lVert\cdot\right\rVert^{q}$ to update $x_{t}^{ag}$.
[…]
Example 3.1 ($p$-norms).
Before presenting our algorithm and its convergence guarantees, we revisit the important example of $p$-norms $\left\lVert\cdot\right\rVert_{p}$, which are regular with $C=1$ by Lemma 2.1. As our distance-generating function, we choose $\psi(x)=\frac{1}{q}\|x\|_{p}^{q}$. The suitable order $q$ depends on $p$. For $p\in(1,2]$, we set $q=2$. This choice ensures that $\psi$ is strongly convex with respect to $\left\lVert\cdot\right\rVert_{p}$ with a rate $\mu=p-1$ (see [4, 21]). For $p>2$, we set $q=p$. This ensures that $\psi$ is $(2^{-\frac{p(p-2)}{p-1}},p)$-uniformly convex with respect to  $\|\cdot\|_{p}$ (see [34, 10]). For both settings, we will show sharp convergence guarantees of our algorithm, as summarized in Table 1.
Example 3.2 (Composite norms).
For composite norms, we can compose distance-generating functions to design new ones. For instance, reconsider the composite norm (2). We know that $\psi_{1}(x_{1:d/2})=\frac{1}{2}\sum_{i=1}^{d/2}x_{i}^{2}$ is $1$-strongly convex with respect to $\|\cdot\|_{2}$ and $\psi_{2}(x_{d/2+1,d})=\frac{1}{1.5}(\sum_{i=d/2+1}^{d}|x_{i}|^{1.5})^{2/1.5}$ is $0.5$-strongly convex with respect to $\|\cdot\|_{1.5}$. Therefore we consider
$\psi(x):=0.5\psi_{1}(x_{1:d/2})+0.5\psi_{2}(x_{d/2+1,d}).$
For all $x,y\in\mathbb{R}^{d}$, it holds
$D_{\psi}(x,y)$ $=0.5\Big{(}\psi_{1}(x_{1:d/2})-\psi_{1}(y_{1:d/2})-\langle\nabla\psi_{1}(y_{1:d/2}),x_{1:d/2}-y_{1:d/2}\rangle\Big{)}$
$+0.5\Big{(}\psi_{2}(x_{(d/2+1):d})-\psi_{2}(y_{(d/2+1):d})-\langle\nabla\psi_{2}(y_{(d/2+1):d}),x_{(d/2+1):d}-y_{(d/2+1):d}\rangle\Big{)}$
$D_{\psi}(x,y)$ $\geq\frac{0.5}{2}\|x_{1:d/2}-y_{1:d/2}\|_{2}^{2}+\frac{0.5\cdot 0.5}{2}\|x_{(d/2+1):d}-y_{(d/2+1):d}\|_{1.5}^{2}\geq 0.5\frac{\|x-y\|_{2\circ 1.5}^{2}}{2},$
therefore $\psi$ is $0.5$-strongly convex with respect to the composite norm $\|\|_{2\circ 1.5}$. In general, the new strongly convex coefficient of the composite norm is the weakest of the two.
Algorithm 1 Accelerated Mirror Descent with Binary Search (for the setting of Assumption 1)
Starting point $x_{1}\in\mathbb{R}$, iterations $T\geq 0$, parameters $(C_{t},\epsilon_{t},\alpha_{t},\eta_{t})_{t\geq 1}$
Set $x_{1}^{ag}=x_{1}$
for $1\leq t\leq T$ do
     Binary search (Algorithm 2) with endpoints $(x_{t},x_{t}^{ag})$ and parameters $(C_{t},\epsilon_{t})$ to find $\lambda_{t}\in[0,1]$, $x_{t}^{md}=\lambda_{t}x_{t}^{ag}+(1-\lambda_{t})x_{t}$ such that $\langle\nabla F(x_{t}^{md}),x_{t}^{md}-x_{t}\rangle+C_{t}\Big{(}F(x_{t}^{md})-F(x_{t}^{ag})\Big{)}\leq\epsilon_{t}$
     Compute $x_{t+1}=\arg\min_{x\in\mathbb{R}^{d}}\{\eta_{t}\langle\nabla F(x_{t}^{md}),x\rangle+D^{\psi}(x,x_{t})\}$
     Compute $x_{t+1}^{ag}=\arg\min_{x\in\mathbb{R}^{d}}\{\alpha_{t}\langle\nabla F(x_{t}^{md}),x\rangle+\frac{\mu}{q}\|x-x_{t}^{md}\|^{q}\}$
end for
return $x_{t+1}^{ag}$
The algorithm takes as input the parameters $(C_{t},\epsilon_{t},\alpha_{t},\eta_{t})$, which include the step sizes $(\alpha_{t},\eta_{t})$ and the stopping conditions $(C_{t},\epsilon_{t})$ for the binary search. All parameters are polynomial functions of the horizon $t$, and are detailed in the proofs in Appendix C.
As we will see in Theorem 4.1 the total number of values and gradients queried in each binary search step is $\mathcal{O}(\log T)$. In turn, the total number of gradients queried in $T$ rounds is $\mathcal{O}(T\log T)$.
3.1 Convergence result
In this subsection, we provide the convergence result of our algorithm. The proof is in Appendix C with a proof sketch in Section 3.2.
Theorem 3.1.
In the setting of Assumption 1, Algorithm 1 with the tuning below returns $x_{T+1}^{ag}$ in $T$ iterations with precision
$F(x_{T+1}^{ag})-F_{\star}~{}=~{}\mathcal{O}_{q,\kappa}\left(\frac{LC^{q}\tau^{q}}{T^{\frac{\kappa q+\kappa-q}{q}}}\Big{(}\frac{D^{\psi}(x_{\star},x_{1})}{\alpha L}+C^{q_{\ast}}\left(\frac{\alpha L}{\mu}\right)^{\frac{\kappa}{q-\kappa}}\log(T)\Big{)}\right)$
where $\mathcal{O}_{q,\kappa}$ omits absolute constants depending only on $(q,\kappa)$. This is achieved for any $\alpha>0$ by
$\alpha_{t}&:=\frac{\alpha}{t^{\beta}}\\ \eta_{t}&:=\alpha_{t}\Big{(}\frac{q-\beta}{\tau C^{q_{\ast}}}t\Big{)}^{q-1}\\ C_{t}&:=C^{q_{\ast}}\Big{(}\frac{\eta_{t}}{\alpha_{t}}\Big{)}^{q_{\ast}-1}-\frac{1}{\tau}=\frac{1}{\tau}\Big{(}(q-\beta)t-1\Big{)}\\ \epsilon_{t}&:=\frac{\alpha^{\frac{q}{q-\kappa}}C^{q_{\ast}}}{t\eta_{t}}\frac{M^{1/r}}{\mu^{1/r}}L^{1+1/r}.\end{aligned}\qquad\text{where}\qquad\begin{aligned} \begin{cases}r&:=\frac{q-\kappa}{\kappa}>0\\ M&:=\big{(}\frac{r}{q}\big{)}^{r}\geq 0\\ \beta&=q-\kappa+\frac{q-\kappa}{q}\end{cases}\end{aligned}$
In particular, we obtain the classic Nesterov accelerated $T^{-2}$ rate for $\kappa=q=2$ and $\alpha\leq\mu/L$.
The optimal tuning for $\alpha$ in the preceding result depends on the precise value of $D^{\psi}(x_{\star},x_{1})$, which is unknown a priori. It also depends on $\log T$, while we would rather keep the algorithm any-time. To avoid both issues, we suggest the following practical tuning.
Corollary 3.2.
Suppose a bound $\frac{1}{\mu}D^{\psi}(x_{\star},x_{1})\leq B$ is available. Then the tuning of Theorem 3.1
[…]
Tuning the parameters. The general step-sizes $\alpha_{t}$ and $\eta_{t}$ are polynomial in the iteration number $t$. The main constraint on them is that their exponent needs to be bounded. We choose $\alpha_{t}=\alpha t^{-\beta}$ with $\alpha>0$ a coefficient we set depending on the information of the Bregman divergence $D^{\psi}(x_{\star},x_{1})$. We further have $\eta_{t}\sim t^{1-\frac{3q-\kappa(1+q)}{q}}$ which gives the classic Lipschitz convex setting $\eta_{t}\sim 1/\sqrt{t}$ for $q=2,\kappa=1$ and the Nesterov step size of $\eta_{t}\sim t$ for $q=\kappa=2$. The choice of $C_{t}\sim t$ is dictated by the mirror descent analysis, and $\epsilon_{t}$ is chosen to decrease sufficiently fast such that $\eta_{t}\epsilon_{t}\lesssim 1/t$ sums to $\mathcal{O}(\log T)$. Full details are provided in Appendix C.
Applying Corollary 3.2 to the $p$-norms of Example 3.1 yields the following convergence rates. We consider a problem where both initialisation $x_{1}$ and minimizer $x_{\star}$ are in the $p$-norm ball of radius $R$, i.e.,  $\left\lVert x_{1}\right\rVert_{p},\left\lVert x_{\star}\right\rVert_{p}\leq R$ (so that $\left\lVert x_{1}-x_{\star}\right\rVert_{p}\leq 2R$ by the triangle inequality). Recall that for $p$-norms, we use distance-generating function $\psi(\cdot)=\frac{1}{q}\left\lVert\cdot\right\rVert_{p}^{q}$ with $q=\max\{2,p\}$. We show in Appendix F that also $D^{\psi}(x_{\star},x_{1})\leq 2R^{q}$.
Corollary 3.3.
In the setting of Assumption 1, with the $p$-norm $\left\lVert\cdot\right\rVert_{p}$ for $1<p\leq 2$, Algorithm 1 with $q=2$ returns an $\epsilon$-optimal solution within
$\mathcal{O}_{\kappa}\left(\left(\frac{L\tau^{2}R^{\kappa}}{\epsilon}\right)^{\frac{2}{3\kappa-2}}\log\Big{(}\frac{L\tau^{\kappa}R^{\kappa}}{\epsilon}\Big{)}\right)$
evaluations of the function value and gradient.
Corollary 3.4.
In the setting of Assumption 1, with the $p$-norm $\left\lVert\cdot\right\rVert_{p}$ for $p\geq 2$, Algorithm 1 with $q=p$ returns an $\epsilon$-optimal solution within
$\mathcal{O}_{p,\kappa}\left(\Big{(}\frac{L\tau^{p}R^{\kappa}}{\epsilon}\Big{)}^{\frac{p}{\kappa p+\kappa-p}}\log\Big{(}\frac{L\tau^{\kappa}R^{\kappa}}{\epsilon}\Big{)}\right)$
evaluations of the function values and the gradients.
These two corollaries for star-convex functions match the lower bounds proved for convex functions, completing Table 1.
3.2 Proof Sketch
The start of the proof is to use star-convexity to linearize the objective $F$ around the gradient point $x_{t}^{md}$, and to relate its performance to that of the mirror descent (be-the-regularized leader) iterates
$x_{t+1}=\arg\min_{x\in\mathbb{R}^{d}}\{\eta_{t}\langle\nabla F(x_{t}^{md}),x\rangle+D^{\psi}(x,x_{t})\}.$
That is, using the star-convexity of $F$, we obtain an upper bound for $F(x_{t}^{md})-F_{\star}$ (see (11)). Since the gradient is evaluated at $x_{t}^{md}$, it is natural to have first a potential evaluated at $x_{t}^{md}$
$\frac{\eta_{t}}{\tau}\Big{(}F(x_{t}^{md})-F_{\star}\Big{)}{}\leq{}$ $D^{\psi}(x_{\star},x_{t})-D^{\psi}(x_{\star},x_{t+1})+\eta_{t}\langle\nabla F(x_{t}^{md}),x_{t}^{md}-x_{t}\rangle$ (3)
[…]
As a next step, we consider the acceleration step, i.e.,
$x_{t+1}^{ag}=\arg\min_{x\in\mathbb{R}^{d}}\{\alpha_{t}\langle\nabla F(x_{t}^{md}),x\rangle+\frac{\mu}{q}\|x-x_{t}^{md}\|^{q}\}.$
After computing the new accelerated point $x_{t+1}^{ag}$ from the proximal step, we transform the previous objective potential analysis of $F(x_{t}^{md})-F_{\star}$ into an accelerated descent lemma comparing the original potential $F(x_{t}^{md})$ with the two neighbour accelerated potentials $\left(F(x_{t}^{ag}),F(x_{t+1}^{ag})\right)$. The key (see (16)) is to combine weak smoothness of $F$ with the inexact gradient-trick [8, 10] (Lemma B.3)
$A_{t}\Big{(}F(x_{t+1}^{ag})-F_{\star}\Big{)}$ $\leq D^{\psi}(x_{\star},x_{t})-D^{\psi}(x_{\star},x_{t+1})+\eta_{t}\langle\nabla F(x_{t}^{md}),x_{t}^{md}-x_{t}\rangle$ (4)
$+$ $\Big{(}A_{t}-\frac{\eta_{t}}{\tau}\Big{)}\Big{(}F(x_{t}^{md})-F(x_{t}^{ag})\Big{)}+\Big{(}A_{t}-\frac{\eta_{t}}{\tau}\Big{)}\Big{(}F(x_{t}^{ag})-F_{\star}\Big{)}+B_{t},$
with $A_{t},B_{t}$ parameters given in (15). The choice of $\lambda_{t}$ found by our binary search Algorithm 2 neutralizes part of that equation, delivering
$\eta_{t}\langle\nabla F(x_{t}^{md}),x_{t}^{md}-x_{t}\rangle+\Big{(}A_{t}-\frac{\eta_{t}}{\tau}\Big{)}\Big{(}F(x_{t}^{md})-F(x_{t}^{ag})\Big{)}\leq\eta_{t}\epsilon_{t}.$
In the end, we obtain the result for iteration $t$
$A_{t}\Big{(}F(x_{t+1}^{ag})-F_{\star}\Big{)}\leq D^{\psi}(x_{\star},x_{t})-D^{\psi}(x_{\star},x_{t+1})+A_{t-1}\Big{(}F(x_{t}^{ag})-F_{\star}\Big{)}+\eta_{t}\epsilon_{t}+B_{t}.$ (5)
For an arbitrary $\alpha>0$, we show that considering $\alpha_{t}=\frac{\alpha}{t^{q-\kappa+\frac{q-\kappa}{q}}}$ leads to a sharp upper bound. We also verify that the choice of parameters satisfies $A_{t}-A_{t-1}\leq\frac{\eta_{t}}{\tau}$, $\eta_{t}\epsilon_{t}=\frac{\mu}{t}$ and $B_{t}\approx\frac{1}{t}$. To conclude, we apply the telescopic sum.
4 Binary search for Hölder smooth functions
In our algorithm (Algorithm 1) we perform binary search to compute the momentum parameter $\lambda_{t}$ that satisfies
$\lambda g^{\prime}(\lambda_{t})+C_{t}g(\lambda_{t})\leq\epsilon_{t},$ (6)
where we use the abbreviation $g(\lambda)\triangleq F(\lambda x_{t}^{ag}+(1-\lambda)x_{t})-F(x_{t}^{ag})$. Condition (6) is later used in the convergence proof of Theorem 3.1. The core idea behind our Algorithm 2 is inspired by Algorithm 2 in [18]. The goal is to find a $\lambda_{t}$ that satisfies (6).
If $g(0)\leq g(1)$ or $g$ is decreasing at $\lambda=1$, then (6) is immediately satisfied. If neither of these conditions hold, then $g(0)>g(1)$ and $g^{\prime}(1)>0$, indicating that $g$ must switch from decreasing to increasing at some $\lambda_{*}\in(0,1)$, thus fulfilling $g^{\prime}(\lambda_{*})=0$.
We further show in the proof of Theorem 4.1 that all $\lambda\in[\lambda_{*}-\delta,\lambda_{*}]$ with a properly chosen $\delta$ also satisfy condition (6). The efficiency of our Algorithm in finding such a $\lambda$ is stated in the Theorem below.
Algorithm 2 Generalised binary search
Objective function $F$, endpoints $(x_{t},x_{t}^{ag})$, parameters $(C_{t},\epsilon_{t})\geq 0$
Define $g(\lambda)\triangleq F(\lambda x_{t}^{ag}+(1-\lambda)x_{t})-F(x_{t}^{ag})$
if  $g^{\prime}(1)\leq\epsilon_{t}$ then
     return $(\lambda_{t}=1,x_{t}^{md}=x_{t}^{ag})$
else if  $C_{t}g(0)\leq\epsilon_{t}$ then
     return $(\lambda_{t}=0,x_{t}^{md}=x_{t})$
end if
Initialise $[a,b]=[0,1]$
loop
     Set $\lambda_{t}=\frac{a+b}{2}$
     if $g^{\prime}(\lambda_{t})+C_{t}g(\lambda_{t})\leq\epsilon_{t}$ then
         return $(\lambda_{t},x_{t}^{md}=\lambda_{t}x^{ag}_{t}+(1-\lambda_{t})x_{t})$
     end if
     if $g(\lambda_{t})>0$ then
         $a=\lambda_{t}$triangleright iterate with $[\lambda_{t},b]$
     else
         $b=\lambda_{t}$triangleright iterate with $[a,\lambda_{t}]$
     end if
end loop
Theorem 4.1.
Under Assumption 1, particularly where $F$ is $(L,\kappa)$-weakly smooth, and for fixed $(C_{t},\epsilon_{t},x_{t},x_{t}^{ag})$ , the binary search algorithm finds $\lambda$, satisfying
$x_{t}^{md}$ $=\lambda x_{t}^{ag}+(1-\lambda)x_{t}$
$\langle\nabla F(x_{t}^{md}),x_{t}^{md}-x_{t}\rangle$ $+C_{t}\Big{(}F(x_{t}^{md})-F(x_{t}^{ag})\Big{)}\leq\epsilon_{t}$
in $\mathcal{O}_{\kappa}\left(\max\{\log(C_{t}),\log\Big{(}\frac{L\|x_{t}-x_{t}^{ag}\|^{\kappa}}{\epsilon_{t}}\Big{)}\}+1\right)$ iterations.
In [18], the binary search algorithm is analyzed for smooth functions, i.e., specifically with $\kappa=2$ in the weakly smooth sense. Our algorithm accommodates a general $\kappa$, addressing the broader class of weakly smooth functions. Although the running time depends on $\kappa$, it appears only as an exponent in the logarithm. Consequently, this results in a change by a constant factor, thereby generalizing the algorithm in [18].
Proof Sketch..
The proof can be roughly divided into three parts. We first show that there exists a $\lambda_{*}$ satisfying the stronger condition (6) with $\epsilon_{t}=0$. Next, we show that this implies that condition (6) holds for all $\lambda\in[\lambda_{*}-\delta,\lambda_{*}]$, where $\delta$ is properly chosen depending on the weak smoothness parameters $(L,\kappa)$ of $F$ and the parameters $(C_{t},\epsilon_{t},x_{t},x_{t}^{ag})$ of the algorithm. Finally we prove that our algorithm finds a $\lambda$ within this interval within at most $\mathcal{O}\Big{(}\log\Big{(}\frac{LC_{t}\|x_{t}^{ag}-x_{t}^{md}\|^{\kappa}}{\epsilon_{t}}\Big{)}\Big{)}$ iterations. ∎
5 Conclusion and future research
In this work, we introduce a novel class of structured nonconvex functions, namely weakly smooth $\tau$-star-convex functions with respect to a norm satisfying the new regularity assumption introduced in Definition 2.7. We demonstrate that the regularity assumption is satisfied not only for any $p$-norm with $p>1$, but also for compositions of regular norms. Consequently, our results apply to a large subclass of star-convex functions, generalizing the setting of smooth star-convex functions with respect to the Euclidean norm [18].
We develop an accelerated algorithm (Algorithm 1) that efficiently minimizes any weakly smooth function within this broad class, extending the binary search technique from [18]. Our analysis shows that the algorithm achieves a near-optimal accelerated convergence rate for all $p$-norms with $p>1$. Additionally, our algorithm has a complexity dependence on $\tau^{\max\{2,p\}}$ for $p\geq 1$, which has been shown to be optimal in the Euclidean case when $p=2$ [18].
While this work focuses on unconstrained minimization problems, future research could extend our approach to constrained settings. Currently, the proximal step for calculating $x_{t+1}^{ag}$ in Algorithm 1 is valid only for unconstrained problems. The main challenge in constrained settings is evaluating the direction of the constrained proximal step.
Our convergence results rely on the assumption of uniform convexity of the distance-generating function $\psi$. While our analysis provides first insights into norms fulfilling this condition, such as (compositions of) $p$-norms, analysing different other regular norms that meet this criterion provides an opportunity for future research.
Finally, our method operates within a first-order access model, meaning that the algorithm has access to both the objective function value and gradient at the evaluation points. It remains an open problem whether this can be extended to “pure” first-order access models, that rely solely on the gradient evaluations. This can be done for the convex case, but it is not known whether this is possible for star-convex functions. Additionally, exploring the applicability of our method to a stochastic model presents another promising direction. Although stochastic accelerated mirror descent has been previously analyzed in the literature [19], the main challenge lies in adapting the binary search algorithm to the noise.
Appendix A Proofs for preliminaries
Recall the notion of $C$-regularity from Definition 2.7.
Lemma A.1.
For $p>1$, $\|\cdot\|_{p}$ is $1$-regular.
Proof.
We denote by $p_{\ast}$ the Hölder conjuguate of $p$, defined such that $\frac{1}{p}+\frac{1}{p_{\ast}}=1$. Recall that for $x\in\mathbb{R}^{d}$, $\|x\|_{p}:=(\sum_{i}|x_{i}|^{p})^{\frac{1}{p}}$, which corresponds to the Euclidean norm for $p=2$. If $p\neq 2$, we observe that for all $x\neq 0$,
$\phi(x)$ $:=\nabla\left(\frac{\|x\|_{p}^{2}}{2}\right)=\frac{1}{2}\nabla\Big{(}(\sum_{i}|x_{i}|^{p})^{2/p}\Big{)}=\frac{1}{2}\cdot\frac{2}{p\|x\|_{p}^{p-2}}\begin{bmatrix}\vdots\\ px_{i}|x_{i}|^{p-2}\\ \vdots\end{bmatrix}=\frac{1}{\|x\|_{p}^{p-2}}\begin{bmatrix}\vdots\\ x_{i}|x_{i}|^{p-2}\\ \vdots\end{bmatrix}.$
As
$\langle\phi(x),x\rangle=\frac{\|x\|_{p}^{p}}{\|x\|_{p}^{p-2}}=\|x\|_{p}^{2}$
and
$\|\phi(x)\|_{\ast}=\|\phi(x)\|_{p_{\ast}}=\frac{\Big{(}\sum_{i=1}^{d}|x_{i}|^{p_{\ast}(p-1)}\Big{)}^{1/p_{\ast}}}{\|x\|_{p}^{p-2}}=\frac{\|x\|_{p_{\ast}(p-1)}^{\frac{p_{\ast}(p-1)}{p_{\ast}}}}{\|x\|_{p}^{p-2}}=\frac{\|x\|_{p_{\ast}(p-1)}^{p-1}}{\|x\|_{p}^{p-2}}=\|x\|_{p},$
where the last equality follows as $p=p_{\ast}(p-1)$, $\|\cdot\|_{p}$ fulfills Definition 2.7 with $C=1$. ∎
Next, we prove regularity for the matrix norms known as Schatten $p$-norms. Recall that, for $p\in[1,\infty]$, the Schatten $p$-norm is defined for any matrix $A\in\mathbb{R}^{n\times m}$ with singular value decomposition $A=\sum_{i}\sigma_{i}u_{i}v_{i}^{\intercal}$ by the $p$-norm of the singular values, i.e.,
$\left\lVert A\right\rVert_{p}~{}:=~{}\Big{(}\sum_{i}\sigma_{i}^{p}\Big{)}^{1/p}.$
The special cases $p=1$, $p=2$ and $p=\infty$ are called the nuclear norm, the Frobenius norm and the spectral norm, respectively.
Lemma A.2.
For $p>1$, the Schatten $p$-norm is $1$-regular.
Proof.
By using, that
$\left\lVert A\right\rVert_{p}~{}:=~{}\Big{(}\sum_{i}\sigma_{i}^{p}\Big{)}^{1/p}=\Big{|}\mathrm{Tr}\left(\left(\sqrt{AA^{T}}\right)^{p}\right)\Big{|}^{1/p}$
by definition, together with the rules of differentiating matrices from Chapter 2.4 in [29], we find
$\phi(A)~{}=~{}\nabla\left(\frac{\left\lVert A\right\rVert_{p}^{2}}{2}\right)~{}=~{}\left\lVert A\right\rVert_{p}^{2-p}\sum_{i}\sigma_{i}^{p-1}u_{i}v_{i}^{\intercal}.$
Hence the inner product satisfies
$\langle\phi(A),A\rangle~{}=~{}\operatorname{tr}\left(\left\lVert A\right\rVert_{p}^{2-p}\sum_{i}\sigma_{i}^{p-1}v_{i}u_{i}^{\intercal}A\right)~{}=~{}\left\lVert A\right\rVert_{p}^{2-p}\operatorname{tr}\left(\sum_{i}\sigma_{i}^{p}v_{i}v_{i}^{\intercal}\right)~{}=~{}\left\lVert A\right\rVert_{p}^{2},$
while also, with $11/p+1/p_{*}=$ 1, the dual norm satisfies
$\left\lVert\phi_{p}(A)\right\rVert_{p_{*}}~{}=~{}\left\lVert A\right\rVert_{p}^{2-p}\left\lVert\sum_{i}\sigma_{i}^{p-1}u_{i}v_{i}^{\intercal}\right\rVert_{p_{*}}~{}=~{}\left\lVert A\right\rVert_{p}^{2-p+p/p_{*}}~{}=~{}\left\lVert A\right\rVert_{p}$
as required. ∎
Lemma A.3.
Let norm $\|\cdot\|_{a}$ defined on $\mathbb{R}^{d_{a}}$ be $C_{a}$-regular and norm $\|\cdot\|_{b}$ defined on $\mathbb{R}^{d_{b}}$ be $C_{b}$-regular. For $\lambda\in(0,1)$, we define the composite norm $\|\cdot\|_{a\circ b}$ as
$\|x\|_{a\circ b}=\left\|\begin{bmatrix}x_{a}\\ x_{b}\end{bmatrix}\right\|_{a\circ b}:=\sqrt{\lambda\|x_{a}\|_{a}^{2}+(1-\lambda)\|x_{b}\|_{b}^{2}},$
where $x_{a}:=x_{1:d_{a}}$ and $x_{b}:=x_{(d_{a}+1):d}$. Then the composite norm $\|\cdot\|_{a\circ b}$ is $\max\{C_{a},C_{b}\}$-regular.
Proof.
Set $\phi_{k}(y):=\nabla\left(\frac{\|y\|_{k}^{2}}{2}\right)$ and abbreviate $y=\begin{bmatrix}y_{1:d_{a}}\\ y_{(d_{a}+1):d}\end{bmatrix}:=\begin{bmatrix}y_{a}\\ y_{b}\end{bmatrix}$. Then
$\phi_{a\circ b}(x)$ $=\nabla\left(\frac{\|x\|_{a\circ b}}{2}\right)=\frac{1}{2}\nabla\left(\lambda\|x_{a}\|_{a}^{2}+(1-\lambda)\|x_{b}\|_{b}^{2}\right)$
$=\frac{1}{2}\begin{bmatrix}\lambda\nabla(\|x_{a}\|_{a}^{2})\\ (1-\lambda)\nabla(\|x_{b}\|_{b}^{2})\end{bmatrix}=\begin{bmatrix}\lambda\phi_{a}(x_{a})\\ (1-\lambda)\phi_{b}(x_{b})\end{bmatrix}.$
As $\|\cdot\|_{a}$ is $C_{a}$-regular and $\|\cdot\|_{b}$ $C_{b}$-regular,
$\langle\phi_{a\circ b}(x),x\rangle$ $=\lambda\langle\phi(x_{1:d_{a}}),x_{1:d_{a}}\rangle+(1-\lambda)\langle\phi(x_{(d_{a}+1):d}),x_{(d_{a}+1):d}\rangle$
$\geq\lambda\|x_{1:d_{a}}\|_{a}^{2}+(1-\lambda)\|x_{(d_{a}+1):d}\|_{b}^{2}=\|x\|_{a\circ b}^{2}.$
Further it follows, that
$\|\phi(x)\|_{\ast}$ $=\sup_{\|z\|\leq 1}\langle z,\phi(x)\rangle=\sup_{\lambda\|z_{a}\|_{a}^{2}+(1-\lambda)\|z_{b}\|_{b}^{2}\leq 1}\langle\begin{bmatrix}z_{a}\\ z_{b}\end{bmatrix},\begin{bmatrix}\lambda\phi_{a}(x_{a})\\ (1-\lambda)\phi_{b}(x_{b})\end{bmatrix}\rangle$
$=\sup_{\alpha\in[0,1]}\left(\sup_{\lambda\|z_{a}\|_{a}^{2}\leq\alpha}\lambda\langle z_{a},\phi_{a}(x_{a})\rangle+\sup_{(1-\lambda)\|z_{b}\|_{b}^{2}\leq 1-\alpha}(1-\lambda)\langle z_{b},\phi_{b}(x_{b})\rangle\right)$
$=\sup_{\alpha\in[0,1]}\left(\sqrt{\alpha\lambda}\sup_{\|z_{a}\|_{a}^{2}\leq 1}\langle z_{a},\phi_{a}(x_{a})\rangle+\sqrt{(1-\alpha)(1-\lambda)}\sup_{\|z_{b}\|_{b}^{2}\leq 1}\langle z_{b},\phi_{b}(x_{b})\rangle\right)$
$=\sup_{\alpha\in[0,1]}\left(\sqrt{\alpha\lambda}\|\phi_{a}(x_{a})\|_{a,\ast}+\sqrt{(1-\alpha)(1-\lambda)}\|\phi_{b}(x_{b})\|_{b,\ast}\right)$
$\leq\sup_{\alpha\in[0,1]}\left(\sqrt{\alpha\lambda}C_{a}\|x_{a}\|_{a}+\sqrt{(1-\alpha)(1-\lambda)}C_{b}\|x_{b}\|_{b}\right)$
$=\sqrt{\lambda C_{a}^{2}\|x_{a}\|_{a}^{2}+(1-\lambda)C_{b}^{2}\|x_{b}\|_{b}^{2}}$
$\leq\max\{C_{a},C_{b}\}\left\lVert x\right\rVert_{a\circ b}.$
This shows that $\|\cdot\|_{a\circ b}$ is $C$-regular with $C=\max\{C_{a},C_{b}\}$ according to Definition 2.7. ∎
Lemma A.4.
For $s\geq 1$, we define $\phi_{s}(x):=\nabla\Big{(}\frac{\|x\|^{s}}{s}\Big{)}.$
[…]
$\begin{cases}\langle\phi_{s}(x),x\rangle&\geq\|x\|^{s}\\ \|\phi_{s}(x)\|_{\ast}^{\alpha}&\leq C^{\alpha}\|x\|^{\alpha(s-1)}.\end{cases}$
Proof.
Recall that $\phi(x)=\nabla(\frac{\|x\|^{2}}{2})$. Then
$\phi_{s}(x)$ $=\nabla\Big{(}\frac{(\|x\|^{2})^{s/2}}{s}\Big{)}=\frac{1}{s}\cdot\frac{s}{2}(\|x\|^{2})^{\frac{s}{2}-1}\cdot\nabla\Big{(}\|x\|^{2}\Big{)}$ $=(\|x\|^{2})^{\frac{s}{2}-1}\cdot\frac{1}{2}\nabla\Big{(}\|x\|^{2}\Big{)}=\frac{\phi(x)}{\|x\|^{2-s}}.$
As
$\langle\phi_{s}(x),x\rangle$ $=\frac{\langle\phi(x),x\rangle}{\|x\|^{2-s}}\geq\frac{\|x\|^{2}}{\|x\|^{2-s}}=\|x\|^{s}$
$\|\phi_{s}(x)\|_{\ast}^{\alpha}$ $=\frac{\|\phi(x)\|_{\ast}^{\alpha}}{\|x\|^{\alpha(2-s)}}\leq\frac{C^{\alpha}\|x\|^{\alpha}}{\|x\|^{\alpha(2-s)}}=C^{\alpha}\|x\|^{\alpha(s-1)},$
the assertion follows. ∎
Appendix B Useful lemma for convergence proof
Lemma B.1.
(Young’s inequality [33]) Let $a,b\geq 0$ and $p,q\geq 1$ such that $\frac{1}{p}+\frac{1}{q}=1$. Then it holds, that
$ab\leq\frac{a^{p}}{p}+\frac{b^{q}}{q}.$
Lemma B.2.
Let $f$ be a convex function and $\nu$ be convex and continuously differentiable. For
$u^{\star}=\arg\min_{u\in{\cal X}}\{f(u)+D^{\nu}(u,y)\},$
and all $u$, it holds
$f(u^{\star})+D^{\nu}(u^{\star},y)+D^{f}(u,u^{\star})\leq f(u)+D^{\nu}(u,y)-D^{\nu}(u,u^{\star}).$
Proof.
As $u^{*}$ is a minimizer, it fulfills for all $u\in{\cal X}$ the first-order optimality condition
$\langle\nabla f(u^{\star})+\nabla D^{\nu}(u^{\star},y),u-u^{\star}\rangle\geq 0,$ (7)
with the gradient taken with respect to the first entry. Applying the three-points identity [8], yields
$\langle\nabla D^{\nu}(u^{\star},y),u-u^{\star}\rangle=D^{\nu}(u,y)-D^{\nu}(u,u^{\star})-D^{\nu}(u^{\star},y),$
leading to
$f(u)-f(u^{\star})-D^{f}(u,u^{\star})$ $=\langle\nabla f(u^{\star}),u-u^{\star}\rangle\geq D^{\nu}(u,u^{\star})-D^{\nu}(u,y)+D^{\nu}(u^{\star},y).$
Lemma B.3.
Set
$\begin{cases}r&:=\frac{q-\kappa}{\kappa}>0\\ M&:=\big{(}\frac{r}{q}\big{)}^{r}\geq 0.\end{cases}$ (8)
For all $\delta>0,x,y\in\mathcal{X}$
$\frac{\|x-y\|^{\kappa}}{\kappa}\leq\frac{M}{q\delta^{r}}\|x-y\|^{q}+\delta.$ (9)
Proof.
Using Young’s inequality B.1 with $\frac{1}{a}+\frac{1}{b}=1$, leads to $t\leq\frac{1}{az}t^{a}+\frac{1}{b}z^{b-1}.$ By choosing $t:=\frac{\|x-y\|^{\kappa}}{\kappa}$, $a=\frac{q}{\kappa}$, $b=\frac{q}{q-\kappa}$, $z=(b\delta)^{\frac{1}{b-1}}=(b\delta)^{r}$, it follows
$\frac{\|x-y\|^{\kappa}}{\kappa}\leq\frac{M}{q\delta^{r}}\|x-y\|^{q}+\delta,$
where we use, that
$\frac{\kappa}{q}\left(\frac{q-\kappa}{q\delta}\right)^{r}\left(\frac{1}{\kappa}\right)^{q/\kappa}=\frac{1}{q\delta^{r}}\left(\frac{q-\kappa}{q\kappa}\right)^{r}=\frac{M}{q\delta^{r}}.$
Appendix C Convergence proof of Theorem 3.1
Our proof can be decomposed in three major steps. We first derive a bound for $F(x_{t}^{md})-F_{\star}$ by analysing our mirror descent step, then we proceed an acceleration step comparing $F(x_{t+1}^{ag})$ and $F(x_{t}^{md})$ and last we derive the parameters satisfying the different conditions needed.
C.1 Mirror descent analysis
We apply Lemma B.2 to the linear function $f(x)=\eta_{t}\langle\nabla F(x_{t}^{md}),x\rangle$ and set $\nu=\psi$, $u^{\star}=x_{t+1}$, $y=x_{t}$, $u=x$. For all $x\in\mathbb{R}^{d}$, this yields
$\eta_{t}\langle\nabla F(x_{t}^{md}),x_{t+1}\rangle+D^{\psi}(x_{t+1},x_{t})\leq\eta_{t}\langle\nabla F(x_{t}^{md}),x\rangle+D^{\psi}(x,x_{t})-D^{\psi}(x,x_{t+1}).$
By setting $x=x_{\star}$ and subtracting $\eta_{t}\langle\nabla F(x_{t}^{md}),x_{\star}\rangle$ on both sides, we obtain
$\eta_{t}\langle\nabla F(x_{t}^{md}),x_{t+1}-x_{\star}\rangle+D^{\psi}(x_{t+1},x_{t})\leq D^{\psi}(x_{\star},x_{t})-D^{\psi}(x_{\star},x_{t+1}).$
Decomposing $\eta_{t}\langle\nabla F(x_{t}^{md}),x_{t+1}-x_{\star}\rangle$ in
$\eta_{t}\langle\nabla F(x_{t}^{md}),x_{t+1}-x_{\star}\rangle=\eta_{t}\langle\nabla F(x_{t}^{md}),x_{t}^{md}-x_{\star}\rangle+\eta_{t}\langle\nabla F(x_{t}^{md}),x_{t+1}-x_{t}^{md}\rangle,$
along with
$\frac{\eta_{t}}{\tau}\Big{(}F(x_{t}^{md})-F_{\star}\Big{)}\leq\eta_{t}\langle\nabla F(x_{t}^{md}),x_{t}^{md}-x_{\star}\rangle,$
following from the the star-convexity of $F$, gives
$\frac{\eta_{t}}{\tau}\Big{(}F(x_{t}^{md})-F_{\star}\Big{)}+\eta_{t}\langle\nabla F(x_{t}^{md}),x_{t+1}-x_{t}^{md}\rangle+D^{\psi}(x_{t+1},x_{t})\leq D^{\psi}(x_{\star},x_{t})-D^{\psi}(x_{\star},x_{t+1}).$
This leads to
$\frac{\eta_{t}}{\tau}\Big{(}F(x_{t}^{md})-F_{\star}\Big{)}+D^{\psi}(x_{t+1},x_{t})$ (10)
[…]
$D^{\psi}(x_{t+1},x_{t})\geq\frac{\mu}{q}\|x_{t+1}-x_{t}\|^{q},$
our equation becomes
$\frac{\eta_{t}}{\tau}\Big{(}F(x_{t}^{md})-F_{\star}\Big{)}$ (11)
[…]
${}+{}\eta_{t}\langle\nabla F(x_{t}^{md}),x_{t}-x_{t+1}\rangle-\frac{\mu}{q}\|x_{t+1}-x_{t}\|^{q}.$
The goal of the next step is to find an upper bound for
$\eta_{t}\langle\nabla F(x_{t}^{md}),x_{t}-x_{t+1}\rangle-\frac{\mu}{q}\|x_{t+1}-x_{t}\|^{q}.$
C.2 Choice of aggregated point
Recall from Algorithm 1 that
$x_{t+1}^{ag}$ $=\arg\min_{x\in\mathbb{R}^{d}}\{\alpha_{t}\langle\nabla F(x_{t}^{md}),x\rangle+\frac{\mu}{q}\|x-x_{t}^{md}\|^{q}\}$
and define $\phi_{q}(x):=\nabla\Big{(}\frac{\|x\|^{q}}{q}\Big{)}$. For all $u\in\mathbb{R}^{d}$, it holds
$\langle\alpha_{t}\nabla F(x_{t}^{md})+\mu\phi_{q}(x_{t+1}^{ag}-x_{t}^{md}),u-x_{t+1}^{ag}\rangle\geq 0.$
Choosing $u=x_{t+1}^{ag}+\lambda(x_{t+1}-x_{t})$ for any $\lambda>0$, leads to
$\lambda\alpha_{t}\langle\nabla F(x_{t}^{md}),x_{t}-x_{t+1}\rangle$ $\leq\lambda\mu\langle\phi_{q}(x_{t+1}^{ag}-x_{t}^{md}),x_{t+1}-x_{t}\rangle.$
Subtracting $\frac{\mu}{q}\|x_{t+1}-x_{t}\|^{q}$ on both sides, yields
$\eta_{t}\langle\nabla F(x_{t}^{md}),x_{t}-x_{t+1}\rangle-\frac{\mu}{q}\|x_{t+1}-x_{t}\|^{q}$
$\leq\frac{\mu\eta_{t}}{\alpha_{t}}\langle\phi_{q}(x_{t+1}^{ag}-x_{t}^{md}),x_{t+1}-x_{t}\rangle-\frac{\mu}{q}\|x_{t+1}-x_{t}\|^{q}$
${}={}\mu\Bigl{\langle}\frac{\eta_{t}}{\alpha_{t}}\phi_{q}(x_{t+1}^{ag}-x_{t}^{md}),x_{t+1}-x_{t}\Bigr{\rangle}-\frac{\mu}{q}\|x_{t+1}-x_{t}\|^{q}$ (12)
Let $q_{\ast}$ be the convex conjugate of $q$, i.e.,  $\frac{1}{q}+\frac{1}{q_{\ast}}=1$. A reformulation of Young’s inequality in Lemma B.1 gives
$\langle a,b\rangle\leq|\langle a,b\rangle|\leq\|a\|_{\ast}\cdot\|b\|$ $\leq\frac{\|a\|_{\ast}^{q_{\ast}}}{{q_{\ast}}}+\frac{\|b\|^{q}}{q},$
leading to
$\langle a,b\rangle-\frac{\|b\|^{q}}{q}$ $\leq\frac{\|a\|_{\ast}^{q_{\ast}}}{{q_{\ast}}}.$
Therefore (C.2) can be upper bounded by
$\mu\Bigl{\langle}\frac{\eta_{t}}{\alpha_{t}}\phi_{q}(x_{t+1}^{ag}-x_{t}^{md}),x_{t+1}-x_{t}\Bigr{\rangle}-\frac{\mu}{q}\|x_{t+1}-x_{t}\|^{q}$ $\leq{}\frac{\mu}{q_{\ast}}\Bigl{\|}\frac{\eta_{t}}{\alpha_{t}}\phi_{q}(x_{t+1}^{ag}-x_{t}^{md})\Bigr{\|}_{\ast}^{q_{\ast}}$ (13)
${}={}\frac{\mu\eta_{t}^{q_{\ast}}}{q_{\ast}\alpha_{t}^{q_{\ast}}}\|\phi_{q}(x_{t+1}^{ag}-x_{t}^{md})\|_{\ast}^{q_{\ast}}$
${}\leq{}\frac{\mu C^{q_{\ast}}\eta_{t}^{q_{\ast}}}{q_{\ast}\alpha_{t}^{q_{\ast}}}\|x_{t+1}^{ag}-x_{t}^{md}\|^{(q-1)q_{\ast}}$
${}={}\frac{\mu C^{q_{\ast}}\eta_{t}^{q_{\ast}}}{q_{\ast}\alpha_{t}^{q_{\ast}}}\|x_{t+1}^{ag}-x_{t}^{md}\|^{q},$
[…]
By setting $u=x_{t}^{md}$, we obtain from first-order optimality, that
$\alpha_{t}\langle\nabla F(x_{t}^{md}),x_{t}^{md}-x_{t+1}^{ag}\rangle$ $\geq\mu\langle\phi_{q}(x_{t+1}^{ag}-x_{t}^{md}),x_{t+1}^{ag}-x_{t}^{md}\rangle,$
leading to
$\alpha_{t}\Big{(}F(x_{t}^{md})-F(x_{t+1}^{ag})\Big{)}$ $\geq\mu\|x_{t}^{md}-x_{t+1}^{ag}\|^{q}-\alpha_{t}D_{F}(x_{t+1}^{ag},x_{t}^{md})$
$\geq\mu\|x_{t}^{md}-x_{t+1}^{ag}\|^{q}-\frac{L\alpha_{t}}{\kappa}\|x_{t}^{md}-x_{t+1}^{ag}\|^{\kappa},$
where we use that $F$ is weakly smooth for the last inequality. By choosing
$r$ $:=\frac{q-\kappa}{\kappa}>0$
$M$ $:=\Big{(}\frac{r}{q}\Big{)}^{r}\geq 0,$
in Lemma B.3, it follows for all $\delta_{t}>0$
$L\alpha_{t}\frac{\|x_{t}^{md}-x_{t+1}^{ag}\|^{\kappa}}{\kappa}\leq\frac{LM\alpha_{t}}{q\delta_{t}^{r}}\|x_{t}^{md}-x_{t+1}^{ag}\|^{q}+L\alpha_{t}\delta_{t}.$
This leads to
$\alpha_{t}\Big{(}F(x_{t}^{md})-F(x_{t+1}^{ag})\Big{)}\geq\Big{(}\mu-\frac{LM\alpha_{t}}{q\delta_{t}^{r}}\Big{)}\|x_{t}^{md}-x_{t+1}^{ag}\|^{q}-L\alpha_{t}\delta_{t}.$
For convenience, we set $\frac{LM\alpha_{t}}{q\delta_{t}^{r}}=\frac{\mu}{q}$ by choosing $\delta_{t}=\Big{(}\frac{L}{\mu}M\alpha_{t}\Big{)}^{1/r}$ and then get
$\alpha_{t}\Big{(}F(x_{t}^{md})-F(x_{t+1}^{ag})\Big{)}\geq\frac{\mu}{q_{\ast}}\|x_{t}^{md}-x_{t+1}^{ag}\|^{q}-\frac{M^{1/r}}{\mu^{1/r}}\Big{(}L\alpha_{t}\Big{)}^{(r+1)/r}.$
Combining the previous equation with (C.2) and (13), yields
$\eta_{t}\langle\nabla F(x_{t}^{md}),x_{t}-x_{t+1}\rangle-\frac{\mu}{q}\|x_{t+1}-x_{t}\|^{q}$ (14)
$\leq$ $\frac{\mu C^{q_{\ast}}\eta_{t}^{q_{\ast}}}{\alpha_{t}^{q_{\ast}}}\frac{\|x_{t+1}^{ag}-x_{t}^{md}\|^{q}}{q_{\ast}}$
$\leq$ $\frac{C^{q_{\ast}}\eta_{t}^{q_{\ast}}}{\alpha_{t}^{q_{\ast}}}\Bigg{(}\alpha_{t}\Big{(}F(x_{t}^{md})-F(x_{t+1}^{ag})\Big{)}+\frac{M^{1/r}}{\mu^{1/r}}\Big{(}L\alpha_{t}\Big{)}^{(r+1)/r}\Bigg{)}$
$=$ $\frac{C^{q_{\ast}}\eta_{t}^{q_{\ast}}}{\alpha_{t}^{q_{\ast}-1}}\Big{(}F(x_{t}^{md})-F(x_{t+1}^{ag})\Big{)}+\frac{C^{q_{\ast}}\eta_{t}^{q_{\ast}}}{\alpha_{t}^{q_{\ast}}}\frac{M^{1/r}}{\mu^{1/r}}\Big{(}L\alpha_{t}\Big{)}^{(r+1)/r}$
C.3 Simplified equation
To simplify notation, we will abbreviate
$A_{t}$ $:=\frac{C^{q_{\ast}}\eta_{t}^{q_{\ast}}}{\alpha_{t}^{q_{\ast}-1}}$ (15)
$B_{t}$ $:=\frac{C^{q_{\ast}}\eta_{t}^{q_{\ast}}}{\alpha_{t}^{q_{\ast}}}\frac{M^{1/r}}{\mu^{1/r}}\Big{(}L\alpha_{t}\Big{)}^{1+1/r}=\frac{C^{q_{\ast}}\eta_{t}^{q_{\ast}}}{\alpha_{t}^{q_{\ast}-1-1/r}}\frac{M^{1/r}}{\mu^{1/r}}L^{1+1/r}.$
Rearranging (11) and combining it with equation (14) from the previous subsections leads to
$A_{t}\Big{(}F(x_{t+1}^{ag})-F_{\star}\Big{)}$ (16)
$\leq$ $D^{\psi}(x_{\star},x_{t})-D^{\psi}(x_{\star},x_{t+1})+\eta_{t}\langle\nabla F(x_{t}^{md}),x_{t}^{md}-x_{t}\rangle$
$+\Big{(}A_{t}-\frac{\eta_{t}}{\tau}\Big{)}\Big{(}F(x_{t}^{md})-F(x_{t}^{ag})\Big{)}$
$+\Big{(}A_{t}-\frac{\eta_{t}}{\tau}\Big{)}\Big{(}F(x_{t}^{ag})-F_{\star}\Big{)}+B_{t}.$
With the binary search on $\lambda_{t}$, Algorithm 1 achieves
$\eta_{t}\langle\nabla F(x_{t}^{md}),x_{t}^{md}-x_{t}\rangle+\Big{(}A_{t}-\frac{\eta_{t}}{\tau}\Big{)}\Big{(}F(x_{t}^{md})-F(x_{t}^{ag})\Big{)}\leq\eta_{t}\epsilon_{t}$
within $\mathcal{O}(\log(\frac{L}{\epsilon_{t}}))$ iterations of function values and gradients. Thus, (16) becomes
$A_{t}\Big{(}F(x_{t+1}^{ag})-F_{\star}\Big{)}\leq$ $D^{\psi}(x_{\star},x_{t})-D^{\psi}(x_{\star},x_{t+1})+\eta_{t}\epsilon_{t}$ (17)
$+\Big{(}A_{t}-\frac{\eta_{t}}{\tau}\Big{)}\Big{(}F(x_{t}^{ag})-F_{\star}\Big{)}+B_{t}.$
To derive the optimal convergence rate, we need the following conditions to be fulfilled for a fixed $\alpha>0$:
$\begin{cases}A_{t}-\frac{\eta_{t}}{\tau}&\leq A_{t-1}\\ B_{t}&\leq\alpha^{\frac{q}{q-\kappa}}\frac{C^{q_{\ast}}}{t}\frac{M^{1/r}}{\mu^{1/r}}L^{1+1/r}.\end{cases}$
If the previous conditions hold, we can compute the telescopic sum in (17) and get
$A_{T}\Big{(}F(x_{T+1}^{ag})-F_{\star}\Big{)}\leq D^{\psi}(x_{\star},x_{1})+\mathcal{O}\Big{(}\alpha^{\frac{q}{q-\kappa}}\frac{C^{q_{\ast}}L^{1+1/r}}{\mu^{1/r}}\log(T)\Big{)}.$ (18)
To find the optimal step-size, we choose
$\beta=q-\kappa+\frac{q-\kappa}{q},\quad\gamma\geq 0,\quad\Big{(}\frac{\eta_{t}}{\alpha_{t}}\Big{)}^{q_{\star}-1}=\frac{q-\beta}{\tau C^{q_{\ast}}}t^{\gamma(q_{\star}-1)}$
and recall that $1+\frac{1}{r}=1+\frac{\kappa}{q-\kappa}=\frac{q}{q-\kappa}$ and $\frac{q_{\ast}}{q}=\frac{1}{q-1}$. Then it follows
$B_{t}\leq\alpha^{\frac{q}{q-\kappa}}\frac{C^{q_{\ast}}}{t}\frac{M^{1/r}}{\mu^{1/r}}L^{1+1/r},$
leading to
$\frac{C^{q_{\ast}}\eta_{t}^{q_{\ast}}}{\alpha_{t}^{q_{\ast}-1-1/r}}\frac{M^{1/r}}{\mu^{1/r}}L^{1+1/r}\leq\alpha^{\frac{q}{q-\kappa}}\frac{C^{q_{\ast}}}{t}\frac{M^{1/r}}{\mu^{1/r}}L^{1+1/r}$
and thus
$\alpha_{t}^{1+1/r}\leq\frac{\alpha^{\frac{q}{q-\kappa}}}{t^{\gamma q_{\ast}+1}}\Leftrightarrow$ $\quad\alpha_{t}\leq\frac{\alpha}{t^{\frac{\gamma(q-\kappa)}{q-1}+\frac{q-\kappa}{q}}}.$
We further note that
$A_{t}-A_{t-1}=C^{q_{\ast}}\Big{(}\frac{\eta_{t}^{q_{\ast}}}{\alpha_{t}^{q_{\ast}-1}}-\frac{\eta_{t-1}^{q_{\ast}}}{\alpha_{t-1}^{q_{\ast}-1}}\Big{)}=(q-\beta)C^{q_{\ast}}\left(\eta_{t}\Big{(}\frac{\eta_{t}}{\alpha_{t}}\Big{)}^{q_{\star}-1}-\eta_{t-1}\Big{(}\frac{\eta_{t-1}}{\alpha_{t-1}}\Big{)}^{q_{\star}-1}\right)\leq\frac{\eta_{t}}{\tau}$
and therefore
$(q-\beta)\Big{(}t^{\gamma(q_{\ast}-1)}-\frac{\eta_{t-1}}{\eta_{t}}(t-1)^{\gamma(q_{\ast}-1)}\Big{)}\leq 1.$
Intuitively, choosing $\eta_{t}$ as a polynomial will ensure that $\frac{\eta_{t-1}}{\eta_{t}}$ tends to $1$. Fixing the parameters in the following way satisfies our condition
$\gamma$ $:=q-1$
$\alpha_{t}$ $:=\frac{\alpha}{t^{\beta}}$
$\eta_{t}$ $:=\alpha_{t}\Big{(}\frac{q-\beta}{\tau C^{q_{\ast}}}t^{\gamma(q_{\star}-1)}\Big{)}^{\frac{1}{q_{\star}-1}}=\alpha_{t}\Big{(}\frac{q-\beta}{\tau C^{q_{\ast}}}t\Big{)}^{q-1}.$
Note that $\gamma(q_{\ast}-1)=1$. We verify that
$\frac{\eta_{t-1}}{\eta_{t}}=$ $\frac{\alpha_{t-1}}{\alpha_{t}}\Big{(}1-\frac{1}{t}\Big{)}^{q-1}=\Big{(}1-\frac{1}{t}\Big{)}^{q-1-\beta}=\Big{(}\frac{t-1}{t}\Big{)}^{q-1-\beta}$
and
$\frac{A_{t}-A_{t-1}}{\eta_{t}}=$ $\frac{q-\beta}{\tau}\Big{(}t-\frac{\eta_{t-1}}{\eta_{t}}(t-1)\Big{)}$
$=$ $\frac{q-\beta}{\tau}\frac{1}{t^{q-1-\beta}}\Big{(}t^{q-\beta}-(t-1)^{q-\beta}\Big{)}$
[…]
$\leq$ $\frac{1}{\tau}\frac{1}{t^{q-\beta-1}}\cdot t^{q-1-\beta}=\frac{1}{\tau}$
By choosing $\epsilon_{t}:=\frac{B_{t}}{\eta_{t}}$ such that $\eta_{t}\epsilon_{t}\leq B_{t}$, leads to
$A_{t}$ $=\frac{C^{q_{\ast}}\eta_{t}^{q_{\ast}}}{\alpha_{t}^{q_{\ast}-1}}=C^{q_{\ast}}\Big{(}\frac{\eta_{t}}{\alpha_{t}}\Big{)}^{q_{\ast}}\alpha_{t}=C^{q_{\ast}}\Big{(}\frac{q-\beta}{\tau C^{q_{\ast}}}t\Big{)}^{\frac{q_{\star}}{q_{\star}-1}}\alpha_{t}$
$\propto C^{q_{\ast}}\cdot\Big{(}\frac{t}{C^{q_{\ast}}\tau}\Big{)}^{q}\cdot\frac{\alpha}{t^{\beta}}=C^{q_{\ast}(1-q)}\cdot\alpha t^{q-\beta}\cdot\frac{1}{\tau^{q}}$
$A_{t}$ $\propto\frac{\alpha}{\tau^{q}C^{q}}t^{\frac{\kappa q+\kappa-q}{q}}$
and finally shows
$F(x_{T+1}^{ag})-F_{\star}$ $=\mathcal{O}_{q,\kappa}\Big{(}\frac{C^{q}\tau^{q}D^{\psi}(x_{\star},x_{1})}{\alpha T^{\frac{\kappa q+\kappa-q}{q}}}+\frac{\alpha^{\frac{\kappa}{q-\kappa}}C^{q_{\ast}}L^{\frac{q}{q-\kappa}}}{\mu^{\frac{\kappa}{q-\kappa}}}\frac{C^{q}\tau^{q}\log(T)}{T^{\frac{\kappa q+\kappa-q}{q}}}\Big{)},$
which proves Theorem 3.1.
Appendix D Proof for binary search
Lemma D.1.
Let $F:\mathbb{R}^{d}\xrightarrow{}\mathbb{R}$ be a $(L,\kappa)$-weakly smooth function. Define $g:\mathbb{R}\to\mathbb{R}$ for $a,b\in\mathbb{R}^{d}$ by
$g(\lambda)=F(a+\lambda b).$
Then for any $\lambda_{1},\lambda_{2}$, we have
$|g^{\prime}(\lambda_{1})-g^{\prime}(\lambda_{2})|\leq\frac{2L}{\kappa}\|b\|^{\kappa}|\lambda_{1}-\lambda_{2}|^{\kappa}.$
Proof.
We recall that :
$g^{\prime}(\lambda)=\langle\nabla F(a+\lambda b),b\rangle$
and for any $\lambda_{1},\lambda_{2}$, by weakly smoothness:
$-\frac{2L}{\kappa}\|\lambda_{1}b-\lambda_{2}b\|^{\kappa}\leq D_{F}(a+\lambda_{1}b,a+\lambda_{2}b)+D_{F}(a+\lambda_{2}b,a+\lambda_{1}b))\leq\frac{2L}{\kappa}\|\lambda_{1}b-\lambda_{2}b\|^{\kappa}$
$-\frac{2L}{\kappa}\|\lambda_{1}b-\lambda_{2}b\|^{\kappa}\leq\langle\nabla F(a+\lambda_{1}b)-\nabla F(a+\lambda_{2}b),b\rangle\leq\frac{2L}{\kappa}\|\lambda_{1}b-\lambda_{2}b\|^{\kappa}$
By taking the absolute value :
$|g^{\prime}(\lambda_{1})-g^{\prime}(\lambda_{2})|\leq\frac{2L}{\kappa}\|\lambda_{1}b-\lambda_{2}b\|^{\kappa}=\frac{2L}{\kappa}\|b\|^{\kappa}|\lambda_{1}-\lambda_{2}|^{\kappa}$
Before proving the theorem, we can make another quick observation
Lemma D.2.
Consider $a<b$ and $g:\mathbb{R}\xrightarrow{}\mathbb{R}$ a function with continuous derivative. Then
$g(a),g^{\prime}(b)>0\geq g(b)\implies\exists\lambda_{\star}\in[a,b],\quad g(\lambda_{\star})\leq 0,g^{\prime}(\lambda_{\star})=0$
Proof.
For that we consider $\lambda_{\star}=\sup_{x\in[0,1]}\{x,g^{\prime}(x)\leq 0\}$ and we know that $\{x,g^{\prime}(x)\leq 0\}$ is not empty because due to Taylor-Lagrange, as $g^{\prime}$ is continuous, there exists $\lambda\in[0,1]$ such that $g^{\prime}(\lambda)=g(1)-g(0)<-\frac{\epsilon_{t}}{C_{t}}\leq 0$. By continuity of $g$ and $g^{\prime}$, we conclude that $g(\lambda_{0})\leq 0,g^{\prime}(\lambda_{0})=0$. ∎
Now we are ready to prove the main theorem.
Theorem D.3.
For fixed $(C_{t},\epsilon_{t},x_{t},x_{t}^{ag})$, the binary search algorithm find a $\lambda$, such that
$x_{t}^{md}$ $=\lambda x_{t}^{ag}+(1-\lambda)x_{t}$
$\langle\nabla F(x_{t}^{md}),x_{t}^{md}-x_{t}\rangle$ $+C_{t}\Big{(}F(x_{t}^{md})-F(x_{t}^{ag})\Big{)}\leq\epsilon_{t}$ (19)
in $\mathcal{O}_{\kappa}\left(\max\{\log(C_{t}),\log\Big{(}\frac{L\|x_{t}-x_{t}^{ag}\|^{\kappa}}{\epsilon_{t}}\Big{)}\}+1\right)$ iterations.
Proof.
Set
$g(\lambda):=F(x_{t}^{md})-F(x_{t}^{ag})=F(\lambda x_{t}^{ag}+(1-\lambda)x_{t})-F(x_{t}^{ag}).$
We notice that $g(1)=0$ and
$g^{\prime}(\lambda)$ $=\langle\nabla F(x_{t}^{md}),x_{t}^{ag}-x_{t}\rangle$
$\langle\nabla F(x_{t}^{md}),x_{t}^{md}-x_{t}\rangle$ $=\langle\nabla F(x_{t}^{md}),\lambda(x_{t}^{ag}-x_{t})\rangle=\lambda g^{\prime}(\lambda).$
As a first step in our proof, we show that there indeed exists a $\lambda$ satisfying (D.3), i.e., fulfilling
$C_{t}g(\lambda)+\lambda g^{\prime}(\lambda)\leq\epsilon_{t}.$
If $\lambda=0$ and $\lambda=1$ fail to satisfy the inequality, we know that both :
$g^{\prime}(1)$ $>\epsilon_{t}$
$C_{t}g(0)$ $>\epsilon_{t}$
From now, we will assume $g^{\prime}(1),g(0)>0$ for the rest of the proof. We set an index set $\Lambda_{\star}=\{\lambda\in[0,1],g(\lambda)\leq 0,g^{\prime}(\lambda)=0\}$. We notice that
$\lambda\in\Lambda_{\star}\implies C_{t}g(\lambda)+\lambda g^{\prime}(\lambda)\leq 0.$
We show that $\Lambda_{\star}$ is not an empty set. For that we will apply the lemma D.2 with $g(0),g^{\prime}(1)>0=g(1)$, there exists $\lambda_{0}\in[0,1],g(\lambda_{0})\leq 0,g^{\prime}(\lambda_{0})=0$ and $\lambda_{0}\in\Lambda_{\star}$.
Now for any $\lambda_{\star}\in\Lambda_{\star}$, since $g^{\prime}$ is $\alpha$-Holder-smooth according to Lemma D.1, for $\delta>0$,
$\forall\lambda\in[\lambda_{\star}-\delta,\lambda_{\star}]:\quad|g^{\prime}(\lambda)-g^{\prime}(\lambda_{\star})|\leq\frac{2L}{\kappa}\|x_{t}-x_{t}^{ag}\|^{\kappa}\delta^{\kappa}$
$\forall\lambda\in[\lambda_{\star}-\delta,\lambda_{\star}]:\quad g(\lambda)-g(\lambda_{\star})\leq\int_{\lambda_{\star}-\delta}^{\lambda_{\star}}|g^{\prime}(t)|dt\leq\frac{2L}{\kappa}\|x_{t}-x_{t}^{ag}\|^{\kappa}\delta^{\kappa+1}$
Therefore, for all $\lambda\in[\lambda_{\star}-\delta,\lambda_{\star}]$,
$C_{t}g(\lambda)+\lambda g^{\prime}(\lambda)$
$~{}=~{}\underbrace{C_{t}g(\lambda_{\star})+\lambda_{\star}g^{\prime}(\lambda_{\star})}_{\leq 0}+C_{t}(g(\lambda)-g(\lambda_{\star}))+\lambda g^{\prime}(\lambda)-\lambda_{\star}\underbrace{g^{\prime}(\lambda_{\star})}_{=0}$
$~{}\leq~{}C_{t}\frac{2L}{\kappa}\|x_{t}-x_{t}^{ag}\|^{\kappa}\delta^{\kappa+1}+\lambda\frac{2L}{\kappa}\|x_{t}-x_{t}^{ag}\|^{\kappa}\delta^{\kappa}$
$~{}\leq~{}\frac{2L}{\kappa}\|x_{t}-x_{t}^{ag}\|^{\kappa}\delta^{\kappa}(C_{t}\delta+1).$
In order to make the last line smaller than $\epsilon_{t}$, we consider $\delta=\min\{\frac{1}{C_{t}},\Big{(}\frac{\kappa\epsilon_{t}}{4L\|x_{t}-x_{t}^{ag}\|^{\kappa}}\Big{)}^{1/\kappa}\}$.
The last step is to notice that in Algorithm 2, our stop condition is satisfied when
$\Lambda_{\star}\cap[a,b]\neq\emptyset,\delta>b-a,$
since that implies $\exists\lambda_{\star}\in\Lambda_{\star},a\in[\lambda_{\star}-\delta,\lambda_{\star}]\implies C_{t}g(a)+ag(a)\leq\epsilon_{t}$.
Now we verifies that the condition $\Lambda_{\star}\in[a,b]$ is hold for during the iterates. For that, using previous lemma D.2, we only need to verifies the condition on the extremity of the segment. If we initially have :
$g(a),g^{\prime}(b)>0\geq g(b)$
then while iterating with $\frac{a+b}{2}$, we know that :
$g\Big{(}\frac{a+b}{2}\Big{)}>0\implies g\Big{(}\frac{a+b}{2}\Big{)},g^{\prime}(b)>0\geq g(b)$ $\implies\Lambda_{\star}\cap\Big{[}\frac{a+b}{2},b\Big{]}\neq\emptyset$
$g^{\prime}\Big{(}\frac{a+b}{2}\Big{)}>0\geq g\Big{(}\frac{a+b}{2}\Big{)}\implies g(a),g^{\prime}\Big{(}\frac{a+b}{2}\Big{)}>0\geq g\Big{(}\frac{a+b}{2}\Big{)}$ $\implies\Lambda_{\star}\cap\Big{[}a,\frac{a+b}{2}\Big{]}\neq\emptyset$
$0\geq g\Big{(}\frac{a+b}{2}\Big{)},g^{\prime}\Big{(}\frac{a+b}{2}\Big{)}$ $\implies\frac{a+b}{2}\in\Lambda_{\star}$
Once we prove that our iteration is valid, we know the interval starts as $[0,1]$ and shrink the size by half at each iteration until reaching $\delta$. So the maximum iteration is :
$\log(1/\delta)+1=\max\{\log(C_{t}),\log\Big{(}\frac{4L\|x_{t}-x_{t}^{ag}\|^{\kappa}}{\kappa\epsilon_{t}}\Big{)}^{1/\kappa}\}+1$
Appendix E Lower bounds in $p$-norm
For the lower bound, we refer to [9] where the authors consider convex functions that are weakly smooth in a non-Euclidean norm.
Theorem E.1.
[9] Let $1\leq p\leq\infty$, and consider the problem class of unconstrained minimization with objectives in the class $\mathcal{F}_{\mathbb{R}^{d},\|\cdot\|_{p}}(\kappa,L)$, whose minima are attained in $\mathcal{B}_{\|\cdot\|_{p}}(0,R)$. Then, the complexity of achieving additive optimality gap $\epsilon$, for any local oracle, is bounded below by:
$\left\{\begin{array}[]{lcl}\Omega\left(\left(\frac{LR^{\kappa}}{\epsilon[\ln d]^{\kappa-1}}\right)^{\frac{2}{3\kappa-2}}\right),&&\mbox{ if }1\leq p<2;\\ \Omega\left(\left(\frac{LR^{\kappa}}{\epsilon\min\{p,\ln d\}^{\kappa-1}}\right)^{\frac{p}{\kappa p+\kappa-p}}\right),&&\mbox{ if }2\leq p<\infty;\text{ and,}\end{array}\right.$
The dimension $d$ for the lower bound to hold must be at least as large as the lower bound itself.
We would like to point out that the previous lower bounds use convex functions whereas the upper bound provided in our work focus on star-convex function class.
Appendix F Bregman divergence and radius of domain
In this section we relate the Bregman divergence generated by the $q$-th power of the $p$-norm to the $q$-th power of the $p$-norms of the arguments.
Lemma F.1.
Fix $p>1$, $q>1$ and define $\psi(x):=\frac{1}{q}\left\lVert x\right\rVert_{p}^{q}$. Then
$D^{\psi}(x,y)~{}\leq~{}2\max\big{\{}\left\lVert x\right\rVert_{p}^{q},\left\lVert y\right\rVert_{p}^{q}\big{\}}.$
Proof.
By definition, we have
$D^{\psi}(x,y)~{}=~{}\frac{1}{q}\left\lVert x\right\rVert_{p}^{q}-\frac{1}{q}\left\lVert y\right\rVert_{p}^{q}-\left\lVert y\right\rVert_{p}^{q-p}\sum_{i}(x_{i}-y_{i})|y_{i}|^{p-2}y_{i},$
which we may reorganize to
${}\leq~{}\frac{1}{q}\left\lVert x\right\rVert_{p}^{q}+\frac{q-1}{q}\left\lVert y\right\rVert_{p}^{q}+\left\lVert y\right\rVert_{p}^{q-p}\sum_{i}x_{i}|y_{i}|^{p-1}.$
Seeing that sum as an inner product, by Hölder with $1=\frac{1}{p}+\frac{1}{\frac{p}{p-1}}$, this is bounded above by
${}\leq~{}\frac{1}{q}\left\lVert x\right\rVert_{p}^{q}+\frac{q-1}{q}\left\lVert y\right\rVert_{p}^{q}+\left\lVert x\right\rVert_{p}\left\lVert y\right\rVert_{p}^{q-1}$
and further by convexity of the exponential (i.e. $a^{\theta}b^{1-\theta}=e^{\theta\ln a+(1-\theta)\ln b}\leq\theta a+(1-\theta)b$) by
${}\leq~{}2\left(\frac{1}{q}\left\lVert x\right\rVert_{p}^{q}+\frac{q-1}{q}\left\lVert y\right\rVert_{p}^{q}\right).$
The result follows by observing that a convex combination is bounded by the maximum. ∎